• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 07:56:33 on June 13, 2001

    Entry number 62962

    keyword=Beam Time Accounting
    Hall A

    6/13/2001 0000-0800
    Experiment #: e99117
    Shift Leader: K. McCormick
    Shift Workers: K. Kramer, B. Michaels

    00000100020003000400050006000700Total Time
    Acceptable Beam
    in Use
    1 35 1 40 45 30 0 0 4.5
    Beam Available
    but Not in Use
    0 10 0 0 0 30 1 1 2.66666666666667
    Beam Not Available
    or Unacceptable
    0 15 0 20 15 0 0 0 0.833333333333333
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 50 1 1 1 30 0 0 5.33333333333333
    Planned Config-
    uration Change
    0 10 0 0 0 0 0 0 0.166666666666667
    Unplanned
    Experiment Down
    0 0 0 0 0 30 1 12.5
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 1 1 2
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Ac 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 12-1: Put in the BeO and checked the position on target. Turned on the raster and check it out. Things check out, going to polarized He target. 1-1:10: NMR Measurement 1:30- Started He measurements (run 1534) 2:30- Run 1535 3:00: Stop run 1535. Beam off for 15 min. 3:15 Beam back. Start run 1536. 4:00 Beam off. 4:17: Beam back. Start run 1537. 5:30: Stopped beam to make NMR measurement 5:45-7: Had target motion problems moving from the pick-up coils. We did several target moves and there seems to be some problem. We are going to have an access to investigate.

    Data Entry by: K. McCormick